Use implicit differenciation to obtain dy over dx

X over y+1 = xsquared+3y


1
Expert's answer
2021-05-19T17:16:48-0400

Given "\frac{x}{y}+1 = x^2 + 3y"


Differentiating both sides,

"\frac{d}{dx}( \frac{x}{y}+1) =\frac{d}{dx}( x^2 + 3y)"


"\frac{1}{y} - \frac{x}{y^2}\frac{dy}{dx} = 2x+3\frac{dy}{dx}"


It can be written as,

"\frac{1}{y} - 2x= 3\frac{dy}{dx}+\frac{x}{y^2}\frac{dy}{dx}"


"\frac{1-2xy}{y}= (3+\frac{x}{y^2})\frac{dy}{dx}"

"\frac{1-2xy}{y} = \frac{3y^2+x}{y^2}\frac{dy}{dx}"


"\frac{dy}{dx} = \frac{(1-2xy)y}{3y^2+x}"